OpenStudy (anonymous):

if the sum of a number and 6 is tripled, the result is four less than twice the number. Find the number

OpenStudy (imstuck):

"if the sum of a number and 6" can be written this way: x+6

OpenStudy (imstuck):

"the result is" means you stick in the equals sign: x+6 =

OpenStudy (imstuck):

"four less than twice the number" is this: 2x-4.

OpenStudy (imstuck):

So the whole equation is x+6=2x-4 Now solve that for x. Can you do that?

OpenStudy (anonymous):

5 ?

OpenStudy (aum):

x + 6 = 2x - 4 subtract x from both sides: 6 = x - 4 add 4 to both sides: 10 = x or x = 10.

OpenStudy (anonymous):

-22

OpenStudy (imstuck):

It's x=10. How did you find it to be -22?

OpenStudy (anonymous):

i dont know but when i typed in 10 for my final check on my homework ...it was wrong and the correct answer was -22